OpenStudy (winsteria):

MEDAL AND FAN Sammie has 3 times as many goldfish as Nora. Nora has 4 goldfish. How many goldfish do Sammie and Nora have in all?

OpenStudy (winsteria):

@KamiBug @TheSmartOne

OpenStudy (kamibug):

N = 4 S = 3*4 S + N = ?

OpenStudy (winsteria):

12 @KamiBug

OpenStudy (kamibug):

Sammie has 12 bc 3*4 = 12. :] Now, to find the total, just add Sammie's amount to Nora's amount, 4. 12 + 4 = ?

OpenStudy (winsteria):

16?

OpenStudy (kamibug):

Correct. :)

OpenStudy (winsteria):

Thank you :)